OpenStudy (anonymous):

HELP PLEASE!!! how do i graph the function f(x)=x^2-9. Also Identify which function you used and the key features of your graph. Explain how to find them algebraically.

OpenStudy (ranga):

If you know how f(x) = x^2 looks like then f(x) = x^2 - 9 is just shifting x^2 vertically down by 9 units.

OpenStudy (ranga):

Key features: find the vertex, the x-intercepts, whether the graph opens up or down, what type of function this is (a parabola), etc.

OpenStudy (anonymous):

The key features of the graph are it's roots (intersection with the x-axis) and intersection with the y-axis. You find the roots by setting f(x) = 0, and the y intersection by setting x = 0 in the expression for f(x). Now you have the key features, but to do it more precisely, you could also calculate the points f(x) for x=-5, x = -4, x = -3, ..., x = 5. Then you will have a good sketch of the graph. :)

OpenStudy (anonymous):

Thank you ranga

OpenStudy (ranga):

you are welcome vicky.

OpenStudy (anonymous):

Yes, ranga; you summed it up well.

OpenStudy (ranga):

thank you. Sfunn

OpenStudy (anonymous):

Thank you as well SfunnSneffanS, it really helped

OpenStudy (anonymous):

Glad I could help :)